### v2

# One Language, Many Gaps: Evaluating Dialect Fairness and Robustness of Large Language Models in Reasoning Tasks

Abstract
Language is not monolithic. While benchmarks, including those designed for multiple languages, are often used as proxies to evaluate the performance of Large Language Models (LLMs), they tend to overlook the nuances of within-language variation, and thus fail to model the experience of speakers of non-standard dialects. Focusing on African American Vernacular English (AAVE), we present the first study aimed at objectively assessing the fairness and robustness of LLMs in handling dialects in canonical reasoning tasks, including algorithm, math, logic, and integrated reasoning. We introduce ReDial (Reasoning with Dialect Queries), a benchmark containing 1.2K+ parallel query pairs in Standardized English and AAVE. We hire AAVE speakers, including experts with computer science backgrounds, to rewrite seven popular benchmarks, such as HumanEval and GSM8K. With ReDial, we evaluate widely used LLMs, including GPT, Claude, Llama, Mistral, and the Phi model families. Our findings reveal that almost all of these widely used models show significant brittleness and unfairness to queries in AAVE. Our work establishes a systematic and objective framework for analyzing LLM bias in dialectal queries. Moreover, it highlights how mainstream LLMs provide unfair service to dialect speakers in reasoning tasks, laying a critical foundation for relevant future research. [^1] Code and data can be accessed here.
One Language, Many Gaps: Evaluating Dialect Fairness and Robustness of Large Language Models in Reasoning Tasks
Fangru Lin1thanks: Work done during internship at Microsoft Research. Contact: fangru.lin@ling-phil.ox.ac.uk, shaoguang.mao@microsoft.com, Shaoguang Mao2, Emanuele La Malfa1,3, Valentin Hofmann4,5 Adrian de Wynter6,7, Xun Wang6, Si-Qing Chen6 Michael Wooldridge1,3, Janet B. Pierrehumbert1, Furu Wei2 1University of Oxford, 2Microsoft Research, 3Alan Turing Institute, 4Allen Institute for AI, 5University of Washington, 6Microsoft Corporation, 7University of York
1 Introduction
Over the last few decades, linguistic research has firmly established that language naturally varies along social, geographic, and demographic dimensions (Chambers and Trudgill, 1998). Such dialectal variation is one of the most salient forms of linguistic diversity. Speakers of “non-standard” dialects are known to experience implicit and explicit discrimination in everyday situations, including housing, education, employment, and the criminal justice system (Baugh, 2005; Adger et al., 2014; Rickford and King, 2016; Drożdżowicz and Peled, 2024). As Large Language Models (LLMs) increasingly serve a broad and rapidly expanding user base (Milmo, 2023; La Malfa et al., 2024), it is critical to understand how they interact with diverse linguistic communities.
In this work, we examine LLMs’ dialect robustness and fairness. For robustness, adversarial robustness provides a consolidated framework to test LLMs on slight variations of existing tasks (Moradi and Samwald, 2021; Jin et al., 2023). Dialects reformulate a problem while maintaining its semantics, i.e., they test what has been referred to as semantic robustness (Malfa and Kwiatkowska, 2022). For fairness, recent research has demonstrated that LLMs exhibit biases against non-standard dialect queries, predominantly assessed in language and social analysis tasks (Sap et al., 2019; Ziems et al., 2023; Hofmann et al., 2024). Equally relevant, yet less studied, are tasks that require reasoning abilities for problem-solving, decision-making, and critical thinking (Wason, 1972; Huth, 2004; Huang and Chang, 2022; Qiao et al., 2022). For instance, algorithm-related tasks (e.g., generation, debugging, etc.) figure prominently in real user queries, as reflected by their first place on the ArenaHard quality board (Li et al., 2024) and their third place on the WildChat frequency board (Zhao et al., 2024).
Figure 1: ReDial is a dialect reasoning benchmark composed of $1,200$+ SE-AAVE parallel queries. Its source data comes from existing benchmarks in SE. AAVE speakers are hired to rewrite each instance in their dialect but preserve their original intent, meaning, and ground truth output label to form their AAVE counterparts.
However, existing dialectal benchmarks (e.g., Ziems et al., 2023) do not cover these tasks, and current popular reasoning benchmarks such as HumanEval (Chen et al., 2021) and GSM8K (Cobbe et al., 2021) are constructed in Standardized English (SE). This could disadvantage dialect speakers in real-world applications like educational assessment (González-Calatayud et al., 2021), personalized recommendation (Kantharuban et al., 2024), and even multimodal tasks (e.g., voice assistant) Martin and Wright (2022), ultimately forcing them to shift their language styles to SE (Cunningham et al., 2024) in order to access the full benefits of modern technologies.
We present the first study that systematically and objectively evaluates LLM fairness and robustness in reasoning tasks expressed in a non-standard dialect. We choose AAVE since around $33$ million people worldwide and approximately $80\%$ of African Americans in the United States speak AAVE, with reports of discriminative behaviors in various scenarios (Lippi-Green, 1997; Purnell et al., 1999; Massey and Lundy, 2001; Grogger, 2011; Rickford and King, 2016). We hire AAVE speakers to manually rewrite instances from seven popular SE reasoning benchmarks into AAVE (Section 2.1). Our approach has unique advantages compared with prior works which either (i) rely on predefined lexical or morphosyntactic transformation rules (Ziems et al., 2022, 2023), which may overlook subtle contextual nuances, or (ii) use LLM as translators (Gupta et al., 2024), which may have the very biases that our research wants to unveil (Fleisig et al., 2024; Smith et al., 2024).
Category Algorithm (25.7%) Logic (29.8%) Math (24.7%) Intergrated (19.7%) Total
Source HumanEval MBPP LogicBench Folio GSM8K SVAMP AsyncHow -
Size 164 150 200 162 150 150 240 1,216
Table 1: ReDial contains $1,216$ fully-annotated parallel prompts for four categories, drawn from seven data sources. Each category contribution to the total amount is reported in percentage points in brackets.
We introduce ReDial (Reasoning with Dialect Queries), the first high-quality, end-to-end human-annotated SE-AAVE parallel dataset for reasoning tasks (Section 2). ReDial contains over 1.2K SE-AAVE prompt pairs covering four canonical reasoning categories: algorithm, math, logic, and integrated reasoning (tasks that require composing multiple reasoning skills). By anchoring these queries to known correct answers and employing human-based rewriting, ReDial furnishes an objective measure of dialect fairness and robustness. It also avoids the pitfalls of LLM-based evaluations, which can be inherently biased (Zheng et al., 2023; Chen et al., 2024; Shi et al., 2024).
Using ReDial, we benchmark widely used LLMs, including GPT-o1, GPT-4o, Claude-3.5-Sonnet, LLaMA-3.1-70B-Instruct, and others (Section 3). We find that almost all models experience statistically significant performance drops on AAVE prompts, despite their semantic equivalence to their SE counterparts. On average, we observe a relative performance reduction of more than 10%. This discrepancy persists even with advanced prompting techniques like Chain-of-Thought (CoT) prompting (Kojima et al., 2022; Wei et al., 2022), indicating that current LLMs are both brittle and unfair with dialectal inputs.
To understand these gaps, we further analyze potential causes. Our analysis reveals that the brittleness of LLMs with AAVE prompts arises from a combination of dialect-specific morphosyntactic features and nuanced conversational norms. Experiments with synthetic perturbations and AAVE-specific feature injections show that while these factors contribute to performance degradation, they fail to replicate the severity observed with human-annotated data. This highlights the limitations of rule-based transformations and the critical need for high-quality, context-rich datasets like ReDial to evaluate LLM fairness and robustness effectively.
In summary, in this paper:
  1. 1.
    We introduce ReDial, the first high-quality, end-to-end human-annotated AAVE-SE parallel reasoning benchmark spanning four foundational reasoning tasks.
  2. 2.
    We show that leading LLMs exhibit significant unfairness and brittleness on AAVE prompts compared to their SE counterparts.
  3. 3.
    We identify that the brittleness of LLMs with AAVE prompts stems from a combination of dialect-specific morphosyntactic features and nuanced conversational norms, which cannot be captured by synthetic transformations.
2 Dataset
ReDial (Reasoning with Dialect Queries) is a benchmark of more than 1.2K parallel Standard English – African American Vernacular English (SE-AAVE) query pairs. Table 1 provides an overview of the distribution, and Figure 1 along with Appendix A.2 present illustrative examples.
Following Zhu et al. (2023a), ReDial includes four canonical reasoning tasks—algorithm, logic, math. We additionally consider integrated reasoning as a compositional task requiring multiple skills. The task formulation of ReDial is linguistically diverse, addresses cornerstone problems in human reasoning, and is of particular interest as it is challenging for LLMs.
We first describe the data sources and sampling strategies (Section 2.1), and then detail the AAVE rewriting and validation processes that ensure high data quality (Section 2.2).
2.1 Data Sourcing
We construct a highly curated dataset by drawing upon seven established benchmarks covering different aspects of reasoning. For each source, we provide key references, task descriptions, and sample sizes. Additional examples can be found in Appendix A.1.
Figure 2: Annotation and cross-validation of ReDial instances. We first sample instances from datasets of four canonical reasoning tasks to compose the source data, then we hire AAVE speakers to rewrite the instances in their dialect. To ensure the quality, we conduct naturalness check by AAVE speakers and correctness check by non-AAVE speakers and LLMs. We reannotate instances that do not pass the quality checks and iterate the process until the data meet our criteria. Finally, we combine the source data and AAVE rewriting to obtain ReDial.
Algorithm HumanEval (Chen et al., 2021) consists of 164 human-written code completion instances. We convert and include all these code completion headings into instruction-following natural language queries following the paradigm of InstructHumanEval. [^2] https://huggingface.co/datasets/codeparrot/instructhumaneval
Algorithm MBPP Austin et al. (2021) includes $1,000$ code generation queries. We randomly sample $150$ instances from its sanitized test set (Liu et al., 2023).
Math GSM8K (Cobbe et al., 2021) is a graduate-level math word problem dataset containing $8,790$ instances. We randomly sample $150$ instances from its test set.
Math SVAMP (Patel et al., 2021) is a collection of $1,000$ elementary-school math problems. We randomly sample $150$ instances from its test set.
Logic LogicBench (Parmar et al., 2024) comprises various logic questions in both binary classification and multiple-choice formats. We sample 100 binary and 100 multiple-choice questions, collecting $200$ samples in total.
Logic Folio (original+perturbed) (Han et al., 2022; Wu et al., 2023) Original Folio is a manually curated logic benchmark. We select $81$ instances along with their manually perturbed versions from Wu et al. (2023), yielding $162$ instances.
Intergrated AsyncHow (Lin et al., 2024) is a planning reasoning benchmark. LLMs must interpret natural language descriptions (i.e., logic), find different possible paths in the graph (i.e., algorithm), and then calculate and compare the time cost for these paths (i.e., math) to reach the correct answer. We use stratified sampling based on the dataset’s complexity metrics and include $240$ instances.
2.2 Annotation and Quality Assurance
After data sourcing, we hire AAVE speakers to rewrite each instance in AAVE. We schematize our annotation and validation in Figure 2 and describe them below.
Annotation.
We hire and ask AAVE speakers to rewrite each SE query so that it sounds natural to AAVE speakers while retaining all critical information (e.g., numerical values, logical conditions, and technical details). For algorithm-related tasks involving code, we hire annotators with a background in computer science to keep the logic and semantics of the code tasks. [^3] Details on annotator compensation, qualifications, and guidelines are presented in Appendix A.3.
Validation.
We then perform a careful quality check to ensure both naturalness and correctness. First, we ask annotators to cross-check and edit each others’ annotations to make sure that the annotations are natural to AAVE speakers. Second, to ensure correctness, we first have non-AAVE speakers manually check the essential information, then conduct a sanity check with GPT-4o for the correctness of rewriting (details in Appendix A.4). We manually check data that GPT-4o flags as invalid to see if all essential information is preserved. No instance is rejected solely based on the LLM’s judgment. We return invalid instances to AAVE speakers for correction and iterate the process until all data pass the check.
After these efforts, we obtain ReDial: a high-quality, end-to-end human-annotated SE–AAVE parallel dataset comprising over 1.2K instances spanning four canonical reasoning tasks. In the rest of this paper, we refer to the SE portion of the dataset as SE ReDial and the AAVE portion as AAVE ReDial.
3 Experiments
3.1 Experimental Setting
3.1.1 Models
We test five families of models, two proprietary and three open-source. The rationale is to benchmark widely used LLMs with impressive reasoning performance. All experiments were conducted between September and December 2024.
GPT.
We use GPT-o1 (OpenAI, ), GPT-4o (OpenAI, 2024), GPT-4 (Achiam et al., 2023), GPT-3.5-turbo (Achiam et al., 2023), [^4] Proprietary model API version information: o1: GPT-o1-preview; gpt-4o: 2024-05-13; gpt-4: 2024-05-03; gpt-3.5-turbo: 2023-11-06. as a family of closed-source models to compare with open-source models for dialect robustness. In particular, o1 is trained using large-scale reinforcement learning (RL) to reason through a chain of thoughts and scales inference time computation to achieve highly complex reasoning paths. It has demonstrated significant improvements in reasoning tasks (OpenAI, ). We use GPT-o1 model to understand how RL reasoning post-training affects LLMs’ dialect robustness and fairness.
Claude.
Developed by Anthropic, the Claude 3 model family represents a widely-used proprietary LLM. For our experiments, we utilize the Claude 3.5 Sonnet model  (Anthropic, ).
LLaMA.
We use LLaMA-3-8B / 70B-Instruct and LLaMA-3.1-70B-instruct (Dubey et al., 2024) which are reported for comparable performance with proprietary GPT models.
Mistral / Mixtral.
We use Mistral-7B-Instruct-v0.3 (Jiang et al., 2023) and Mixtral-8x7B-Instruct-v0.1 (Jiang et al., 2024). Mistral-7B-Instruct-v0.3 is reported to be outstanding in reasoning; with Mixtral-8x7B-Instruct-v0.1, we can understand whether Mixture-of-Expert (MoE) architectures enhance dialect robustness.
Phi.
We use Phi-3-Mini / Small / Medium-128K-Instruct (Abdin et al., 2024; Gunasekar et al., 2023) in our experiment. Phi-3 models, pre-trained on carefully designed “textbook” data, are reported for impressive performance in reasoning despite their small sizes (3.8/7/14B parameters each). We use these models to understand how highly curated pre-training data affect LLMs’ dialect robustness and fairness.
Model Setting Algorithm Logic Math Integrated All
SE AAVE SE AAVE SE AAVE SE AAVE SE AAVE $\Delta$
GPT-o1 Direct 0.818 0.825 0.947 0.923 0.878 0.815 0.942 0.925 0.892 0.866 -0.026
GPT-4o Direct 0.790 0.761 0.933 0.930 0.818 0.768 0.783 0.312 0.832 0.716 -0.116
CoT 0.771 0.761 0.950 0.920 0.815 0.771 0.762 0.662 0.826 0.784 -0.043
GPT-4 Direct 0.742 0.723 0.840 0.713 0.796 0.749 0.217 0.133 0.678 0.612 -0.067
CoT 0.723 0.608 0.920 0.813 0.793 0.743 0.283 0.058 0.706 0.590 -0.115
GPT-3.5-turbo Direct 0.653 0.631 0.667 0.443 0.533 0.544 0.200 0.129 0.531 0.460 -0.072
CoT 0.646 0.551 0.753 0.543 0.503 0.425 0.075 0.067 0.517 0.416 -0.101
Claude-3.5-Sonnet Direct 0.771 0.806 0.970 0.930 0.851 0.776 0.879 0.717 0.865 0.810 -0.055
CoT 0.774 0.736 0.953 0.940 0.859 0.796 0.900 0.771 0.868 0.811 -0.058
LLaMA-3.1-70B Direct 0.726 0.653 0.767 0.893 0.702 0.630 0.392 0.112 0.663 0.599 -0.064
CoT 0.723 0.653 0.880 0.870 0.809 0.768 0.579 0.500 0.759 0.711 -0.049
LLaMA-3-70B Direct 0.682 0.643 0.907 0.887 0.663 0.552 0.158 0.067 0.628 0.562 -0.066
CoT 0.697 0.646 0.923 0.887 0.616 0.561 0.517 0.350 0.693 0.622 -0.072
LLaMA-3-8B Direct 0.535 0.510 0.827 0.800 0.478 0.464 0.025 0.067 0.489 0.480 -0.009
CoT 0.532 0.478 0.827 0.800 0.475 0.492 0.029 0.025 0.488 0.472 -0.016
Mixtral-8x7B Direct 0.452 0.401 0.520 0.340 0.414 0.240 0.100 0.075 0.388 0.274 -0.114
CoT 0.468 0.411 0.687 0.567 0.384 0.285 0.133 0.071 0.431 0.345 -0.086
Mistral-7B Direct 0.331 0.255 0.400 0.213 0.315 0.271 0.096 0.075 0.297 0.214 -0.083
CoT 0.312 0.245 0.453 0.347 0.323 0.293 0.083 0.083 0.305 0.252 -0.053
Phi-3-Medium Direct 0.545 0.433 0.867 0.790 0.500 0.470 0.050 0.038 0.513 0.454 -0.059
CoT 0.548 0.455 0.860 0.827 0.492 0.439 0.067 0.029 0.513 0.458 -0.055
Phi-3-Small Direct 0.615 0.252 0.820 0.760 0.530 0.525 0.058 0.062 0.530 0.421 -0.109
CoT 0.570 0.194 0.893 0.843 0.544 0.522 0.096 0.079 0.549 0.429 -0.119
Phi-3-Mini Direct 0.557 0.427 0.520 0.550 0.605 0.525 0.021 0.042 0.456 0.410 -0.046
CoT 0.576 0.443 0.773 0.750 0.622 0.528 0.017 0.021 0.528 0.461 -0.067
Table 2: We report model pass rates using direct and CoT prompting on ReDial, including individual performances on subtasks and overall performance/gap (in column all). We follow the recommendations from Dror et al. (2018) and test the statistical significance of performance differences between SE and AAVE on all results using the McNemar’s test for binary data (McNemar, 1947). We correct p-values for multiple measurements using the Holm-Bonferroni method (Holm, 1979). Statistically significant drops are in bold. Details in Appendix A.7.
3.1.2 Implementation and Evaluation
Implementation.
We set the temperature to zero for the main experiments to ensure maximum reproducibility. We report two prompting methods in our main results: (i) direct prompting LLMs with task instances, which resembles general real-life use cases the most (Direct) and (ii) zero-shot Chain of Thought (Wei et al., 2022; Kojima et al., 2022) (CoT, i.e., adding instructions in the spirit of “Let’s think step by step” on top of task descriptions, which resembles expert user prompts to improve model performance). For GPT-o1, we only test the direct prompting as its inherent CoT reasoning pattern. [^5] We also test non-zero temperatures and report results in Appendix A.6. We report further implementation details in Appendix A.5.666We deliberately avoid testing advanced prompting methods, such as Tree of Thought (Yao et al., 2024) and Self-Refine (Madaan et al., 2024). Our focus is on evaluating how LLMs perform when prompted for everyday use by dialect users, which is critical for assessing fairness in LLMs. Similarly, we do not fine-tune any models, as our study aims to investigate biases inherent in base models. The effects of fine-tuning are beyond the scope of this study.
Algorithm Math Logic Intergrated All
SE 0.632 0.622 0.768 0.302 0.597
AAVE 0.563 0.564 0.706 0.212 0.529
$\Delta$0.069 -0.058 -0.062 -0.090 -0.068
Table 3: Pass rates by task averaged across responses from all models with direct prompting. In bold, results statistically deviate according to McNemar’s tests applied to AAVE and SE (i.e., models have significant drops in AAVE). We also report the SAE-AAVE absolute delta in performance.
Evaluation.
To unify evaluation metrics, we consider the pass rate for all tasks. For Algorithm, we consider Pass@1 using all base and extra unit test cases in EvalPlus (Liu et al., 2023), which results in either pass or fail for every code generation. We convert all other task measures of correctness or incorrectness to pass or fail.
3.2 Experimental Results
We report pass rates for ReDial in Table 2 and  3, and summarize the main results of our experiments.
All Models are Brittle.
All models experience performance drops in AAVE compared to SE ReDial, and these drops are statistically significant in all cases, with the sole exception of LLaMA-3-8B-Instruct. This indicates that our benchmark poses huge challenges to models, both in terms of absolute performance and with respect to their dialect robustness and fairness.
The absolute performance gaps commonly range from around 5% to over 10% ($\Delta$ in table 2). Specifically, GPT-4o (zero-shot) shows an absolute gap of 11.6%, dropping from an average of 0.832 to 0.716. GPT-4 (CoT) exhibits an 11.5% drop. Mixtral-8x7B-Instruct-v0.1 (zero-shot) shows a particularly large difference of 11.4% points as well. Interestingly, we found that although the performance drop of GPT-o1 is smaller than other GPT models, but still significant. It indicates that although further RL post-training on general reasoning and inference scaling can systematically enhance dialect robustness and fairness, they cannot completely solve the problem.
In short, dialect unfairness and brittleness are identified in all the models we examined, including the MoE and large reasoning models. This finding indicates that the problem is widespread, non-trivial, and cannot be easily mitigated by naively changing model architecture or proposing more complex reasoning paths.
All Tasks are Brittle.
When aggregated by task type, AAVE queries cause a statistically significant performance drop across all these categories (Table 3). For instance, when averaging results across all models: direct prompting leads to an average 10% relative performance drop.
Interestingly, integrated reasoning tasks, which require multiple reasoning skill compositions, show some of the largest relative drops (about 30%). This suggests that compositionally complex task may be more prone to dialect brittleness.
Prompting and Inference Scaling are Brittle.
While CoT prompting can slightly reduce the discrepancy for some models, it fails to close it entirely. For example, GPT-4o’s performance gap decreases from about 0.116 (zero-shot) to 0.055 (CoT). This suggests that even when models are given additional reasoning "scaffolding," their understanding and performance in AAVE remain comparatively weaker than in SE, which is also in line with our observation with GPT-o1 results.
Model Scaling is Brittle.
All model families display some degree of dialect-related performance degradation. A notable observation is that simply using larger models does not inherently improve robustness to AAVE. For example, even LLaMA-3.1-70B-Instruct, among the largest and most capable tested models, suffers from significant performance drops on AAVE queries. This pattern holds across the board, indicating that scaling model size alone is insufficient to address dialect-related performance disparities.
4 Discussions
This section investigates the potential reasons for AAVE’s brittleness. We show that LLMs’ brittleness with AAVE reasoning queries cannot be simply attributed to the lack of understanding of this dialect or simple lexical features. The nuanced conversational norms of AAVE also play an important role in the problems for LLMs.
4.1 General Understanding and Morphosyntactic Features
One possible explanation for the performance drop is that LLMs cannot process AAVE dialect. We thus computed perplexities on ReDial AAVE vs. SAE prompts. Indeed, Table 4 confirms that LLMs exhibit higher perplexities on dialect than SE.
Models SE AAVE
LLaMA-3.1-70B-Instruct 9.4 17.5
Phi-3-Medium-128K-Instruct 5.9 12.5
Phi-3-Mini-128K-Instruct 7.1 15.9
Table 4: Averaged perplexities across instances calculated by different models on SE/AAVE ReDial.
However, is the insufficient understanding the only reason that is causing LLMs’ performance to drop? To answer it, we gradually inject typos in SE ReDial by replacing/deleting/adding words/characters, such that we make the input texts more difficult for LLMs (i.e., the measured perplexity goes up). We find that while these perturbations degrade model performance, the drop does not reach the severity observed with natural AAVE data on large-scale models (see full results in Section A.8). This discrepancy suggests that AAVE brittleness is not solely due to the general difficulty in processing. There are some dialect-specific reasons for the performance drop.
If language-agnostic processing ability cannot explain LLMs’ brittleness, can we attribute the problem to morphosyntactic AAVE features? Following Ziems et al. (2022, 2023), we use morphosyntactic transformation rules to inject AAVE features into SE ReDial. We find that performance degradation generally intensifies as the density of AAVE-specific features increases when we gradually inject AAVE features into standard ReDial (see full results in Section A.9). This suggests that these features play a significant role in diminishing model performance.
However, even under the most extreme synthetic perturbations, performance drops are notably less severe than those observed with human-rewritten prompts. This underscores the critical importance of our high-quality human-annotated dialect data ReDial for evaluating LLM fairness and robustness. Synthetic rule-based transformations provide valuable insights, yet fail to capture the contextual depth of real-world dialect usage.
4.2 AAVE Conversational Norms
We use the mutual information between the token distributions of SE and AAVE ReDial to find that the top 5 most informative AAVE features in terms of distinguishing them from SE are , up, in, gon, and gotta. We note that many of these features are not well-known AAVE-specific features (e.g., up). Through a further investigation of our dataset, we find that these lexicons are associated with phrase-level AAVE constructions. For instance, instead of saying …encode the answer… in Standard English, AAVE instruction says …wrap it up…. This finding is particularly interesting as we find that, in addition to previous linguistic observations of AAVE-specific morphosyntactic features, there are conversational norms of the dialect such as nuanced uses of phrases.
We compute Spearman’s correlation between the frequency of the features we find in each instance and their corresponding performance drop. Indeed, these features play a significant role in predicting GPT-4o’s performance degradation (r=-0.318, p<0.001). We further implement and analyze 12 rule-based AAVE features following Ziems et al. (2022) (details in Appendix A.11), which are well documented in linguistic literature such as "finna" as a maker of immediate future Nguyen and Grieve (2020). We notice that the influential lexical features are a subset of the feature set discovered by mutual information (i.e., some of the actual influential features are not encoded in synthetic transformation rules). Consequently, the influence of synthetic features is not as strong as those discovered by mutual information (r=-0.256, p<0.001). This means that simple rule-based transformations that implement the most salient morphosyntactic AAVE features may not be able to capture rich, context-dependent use of the dialect and, therefore fall short in predicting LLMs’ performance in real workflows.
Aided by GPT-o1 preview to filter the vast amount of data, we conducted a linguistically informed analysis of LLMs’ most frequent errors on AAVE. For the algorithmic tasks, grammatical constructions and non-standard verb forms (e.g., finna’, ’em), omission of articles and auxiliary verbs may cause the model to misinterpret references and function naming conventions. For example, GPT-4o interprets you gon’ write a python function, python_function as a general statement rather than a directive to name the function. On logic tasks, the frequent use of double negatives, zero copula, and inverted conditionals introduces structural ambiguities. For example, the construct He don’t take no breaks can invert intended meanings, leading the model to misunderstand conditional statements. On math, informal expressions, unclear quantity references, and non-standard comparative cause erroneous parsing of numerical information and confusion over collective versus individual quantities. Informal phrasings like half as much as he be runnin’ and ambiguous comparative expressions (4 fewer boxes of apple pie than on Sunday) can cause the model to misinterpret numerical relationships, resulting in errors when calculating totals, differences, or fractions. On the integrated task, phonetic spellings, colloquial connectors, and inverted word orders limit the model’s understanding of concurrency, and stepwise instructions. Such dialectal nuances highlight the necessity of our dataset and also call for more efforts to collect more human data for relevant purposes.
5 Related Work
Dialect studies in natural language processing.
Previous works on AAVE studies in natural language processing mostly focus on non-reasoning-heavy tasks such as POS tagging (Jørgensen et al., 2015, 2016), language identification and dependency parsing (Blodgett et al., 2016), automatic captioning (Tatman, 2017), and general language generation (Deas et al., 2023). AAVE is also found to be more likely to trigger false positives in hate speech identifiers (Davidson et al., 2019; Sap et al., 2019) due to specific word choices (Harris et al., 2022), be considered negative by automatic sentiment classifier (Groenwold et al., 2020), and cause covert biases in essential areas of social justice (Hofmann et al., 2024). Relevant studies (Ziems et al., 2022; Gupta et al., 2024) also find that rule-based AAVE feature perturbations can downgrade language model performance in a wide range of tasks covered by GLUE (Wang, 2018).
More generally, dialects in world languages pose challenges to natural language processing systems. Ziems et al. (2023) find that auto-encoder models are brittle on rule-based English dialect feature perturbations. Fleisig et al. (2024) report that English dialect speakers perceive responses generated by chatbots to be more negative than Standardized English (SE) prompts. Faisal et al. (2024) find that world dialects cause problems in tasks including dependency parsing (Scherrer et al., 2019) and machine translation (Mirzakhalov, 2021) on mBERT and XLM-R (Conneau et al., 2020).
Fairness and Robustness of Large Language Models. LLMs are widely testified to be both unfair and brittle. They introduce unfair performance (Huang et al., 2023; Dong et al., 2024) and cost (Petrov et al., 2024) to users across different languages, exacerbate social imbalance by marginalizing minority groups in various aspects including gender (Kotek et al., 2023; Fraser and Kiritchenko, 2024), race (Hofmann et al., 2024; Wang et al., 2024), and culture (Naous et al., 2023; Tao et al., 2024). They also show different performance to reasoning tasks in different languages (Huang et al. 2023, 2024; Ranaldi et al. 2024; inter alia). For the first time, our study provides extensive empirical evidence that LLMs exhibit unfairness in reasoning tasks. This bias specifically affects speakers of certain dialects within a single language.
Previous works report that LLMs are very sensitive to slight variations of prompts by introducing typos or paraphrasing in SE (Elazar et al., 2021; Liang et al., 2022; Raj et al., 2022; Zhu et al., 2023b; Lin et al., 2024). In this work, we consider a novel application of using human-written perturbations in AAVE by asking humans to rewrite instances to their dialect and evaluate LLM robustness towards these natural perturbations, which have proven to cause LLMs to be more brittle than synthetic typo-style (Section 4.1) or linguistic-rule-based (Appendix A.9) perturbations.
6 Conclusion
Our study is the first attempt to objectively evaluate the dialect robustness and fairness of LLMs across reasoning tasks. We introduce ReDial, a dataset comprising over 1.2K parallel prompts in Standardized English and African American Vernacular English (AAVE) tailored to reasoning tasks: algorithm, logic, math, and integrated reasoning. Extensive empirical evidence on ReDial demonstrates that LLMs exhibit significant unfairness and brittleness when reasoning tasks are expressed in AAVE. These findings underscore the unfairness to dialect users and LLMs’ brittleness with natural prompt variations with the same semantics. We advocate for further research to enhance dialect fairness and robustness of LLMs, ensuring equal service for all linguistic groups and demographics.
7 Limitations
First, as the first systematic framework for analyzing LLM bias in dialectal queries for reasoning tasks, we selected AAVE due to its linguistic significance and cultural impact. However, we recognize the vast diversity of dialects worldwide. The insights derived from AAVE may not generalize to other dialects. To ensure annotation quality and maintain the focus of our study, we concentrated on AAVE with high-quality human annotations. Future research could expand on our framework to encompass a wider range of dialects, generating more broadly applicable conclusions.
Second, our benchmark, ReDial, evaluates LLM performance across four categories of reasoning tasks using queries sampled from seven popular and well-documented benchmarks. While these tasks are representative of common reasoning challenges, we acknowledge that reasoning is a multifaceted domain with many additional categories and tasks that fall outside the scope of this study.
Third, we evaluated five representative LLM families in this study, including widely used and sota models. However, given the rapid proliferation of new LLMs, testing every model is infeasible. We hope that future research will use the ReDial benchmark to investigate fairness and reasoning robustness across a broader range of LLMs as they emerge.
Last but not least, while we present extensive empirical evidence demonstrating the performance drop of LLMs on dialectal queries, our study does not deeply investigate the underlying causes of these performance discrepancies or propose systematic methods to mitigate this bias. These topics exceed the scope of our work but are critical for addressing the inequities we have identified. Despite this limitation, we believe that ReDial provides a robust and systematic tool to help researchers explore these issues. The absence of immediate solutions should not detract from the significance of our findings, which lay the groundwork for future efforts to address fairness and robustness in LLMs.
8 Ethic Statement
ReDial is a collection of high-quality human-annotated translations: obtaining such data requires making clear design choices and poses ethical questions that we hereby address.
For data collection, we deliberately do not set hard constraints for annotator identity and demographic verification, recognizing there are no definite boundaries to identify dialects and their speakers (King, 2020). The authors further elaborate that the term “AAVE” itself is contested, with alternatives that could be used instead; in employing the term “AAVE”, we adhere to the widely used terminology in related works on dialects and NLP (Ziems et al., 2022; Gupta et al., 2024). We corroborate the data quality by asking self-identified dialect speakers to cross-validate each others’ answers.
We do not collect annotators’ personal information; while we firmly commit to this rule to protect annotators’ privacy, it makes it difficult to draw conclusions about how annotators’ backgrounds shape their writing/individual-level variations. Further on the ethical aspect of data collection, we work with a data vendor that makes sure the recruitment and annotation adhere to high standards for and from the annotators. However, although we have a legal contract and we try our best to convey our guidelines and requirements, we admit that we do not have full control over how the vendor recruits people and conducts data annotation.
We also stress that the LLM validation stage in our quality control process is not completely trustworthy as even they are prone to hallucinations (Ji et al., 2023) and biases against minority groups (Xu et al., 2021; Fleisig et al., 2024; Smith et al., 2024; Wang et al., 2024). To mitigate this issue, we conduct full manual checks of every instance identified as invalid by an LLM so that no instance is rejected purely because of LLM decisions.
9 Acknowledgement
We thank all the bodies who have provided funding for the authors and for the associated project. FL is supported by Clarendon and Jason Hu studentship. ELM is supported by the Alan Turing Institute. We thank Ambrosio Blanco for helpful feedback on the ethical review. We are grateful to the people who offered feedback and suggestions along the way, and in particular, Janet B. Pierrehumbert, Su Lin Blodgett, Wenshan Wu, Xiaoyuan Yi, Yan Xia and other colleagues in Microsoft Research, who offered invaluable advice and helped us refine the paper.
Appendix A Appendix
A.1 Source Dataset Illustration
A.1.1 Algorithm
Original HumanEval
from typing import List def has_close_elements(numbers: List[float], threshold: float) -> bool: """Checkifingivenlistof ␣␣␣␣numbers,areanytwonumbers ␣␣␣␣closertoeachotherthan ␣␣␣␣giventhreshold. ␣␣␣␣$>$has_close_elements([1.0, ␣␣␣␣2.0,3.0],0.5) ␣␣␣␣False ␣␣␣␣$>$has_close_elements([1.0, ␣␣␣␣2.8,3.0,4.0,5.0,2.0],0.3) ␣␣␣␣True ␣␣␣␣""" InstructHumanEval Used in the Paper
Write a function has_close_elements(numbers: List[float], threshold: float) -$>$ bool to solve the following problem: Check if in given list of numbers, are any two numbers closer to each other than given threshold. $>$much-greater-thanabsent>>>> > > has_close_elements([1.0, 2.0, 3.0], 0.5) False $>$much-greater-thanabsent>>>> > > has_close_elements([1.0, 2.8, 3.0, 4.0, 5.0, 2.0], 0.3) True
MBPP
Write a python function to remove first and last occurrence of a given character from the string. Your code should pass these tests: assert remove_Occ(“hello",“l") == “heo" assert remove_Occ(“abcda",“a") == “bcd" assert remove_Occ(“PHP",“P") == “H"
A.1.2 Logic
LogicBench
If an individual consumes a significant amount of water, they will experience a state of hydration. Conversely, if excessive amounts of sugar are ingested, a sugar crash will ensue. It is known that at least one of the following statements is true: either the Jane consumes ample water or she will not experience a sugar crash. However, the actual veracity of either statement remains ambiguous, as it could be the case that only the first statement is true, only the second statement is true, or both statements are true. Can we say at least one of the following must always be true? (a) she will feel hydrated and (b) she doesn’t eat too much sugar
Folio
Consider the following premises: “People in this club who perform in school talent shows often attend and are very engaged with school events. People in this club either perform in school talent shows often or are inactive and disinterested community members. People in this club who chaperone high school dances are not students who attend the school. All people in this club who are inactive and disinterested members of their community chaperone high school dances. All young children and teenagers in this club who wish to further their academic careers and educational opportunities are students who attend the school. Bonnie is in this club and she either both attends and is very engaged with school events and is a student who attends the school or is not someone who both attends and is very engaged with school events and is not a student who attends the school." Assuming no other commonsense or world knowledge, is the sentence “Bonnie performs in school talent shows often." necessarily true, necessarily false, or neither? Answer either “necessarily true", “necessarily false", or “neither".
A.1.3 Math
GSM8K
Given a mathematics problem, determine the answer. Simplify your answer as much as possible and encode the final answer in $<$answer><absent><> </answer$>$ (e.g., $<$answer$>$1$<$/answer$>$). Question: Janet’s ducks lay 16 eggs per day. She eats three for breakfast every morning and bakes muffins for her friends every day with four. She sells the remainder at the farmers’ market daily for \$2 per fresh duck egg. How much in dollars does she make every day at the farmers’ market? Answer:
SVAMP
Given a mathematics problem, determine the answer. Simplify your answer as much as possible and encode the final answer in $<$answer><absent><> </answer$>$ (e.g., $<$answer$>$1$<$/answer$>$). Question: Winter is almost here and most animals are migrating to warmer countries. There are 41 bird families living near the mountain. If 35 bird families flew away to asia and 62 bird families flew away to africa How many more bird families flew away to africa than those that flew away to asia? Answer:
A.1.4 Comprehensive
AsyncHow
To create a video game, here are the steps and the times needed for each step. Step 1. Learn the basics of programming (180 days) Step 2. Learn to use a language that is used in games (60 days) Step 3. Learn to use an existing game engine (30 days) Step 4. Program the game (90 days) Step 5. Test the game (30 days)
These ordering constraints need to be obeyed when executing above steps: Before starting step 2, complete step 1. Before starting step 3, complete step 1. Before starting step 4, complete step 2. Before starting step 4, complete step 3. Before starting step 5, complete step 4.
Question: Assume that you need to execute all the steps to complete the task and that infinite resources are available. What is the shortest possible time to create a video game? Answer the time in double quotes. Answer:
A.2 ReDial Samples
Algorithm
Standardized Write a function python_function(numbers: List[float], threshold: float) >->- > bool to realize the following functionality: Check if in given list of numbers, are any two numbers closer to each other than given threshold. $>$much-greater-thanabsent>>>> > > python_function([1.0, 2.0, 3.0], 0.5) False $>$much-greater-thanabsent>>>> > > python_function([1.0, 2.8, 3.0, 4.0, 5.0, 2.0], 0.3) True Generate a Python function to solve this problem. Ensure the generated function is named as python_function.
Algorithm
AAVE Aight, so here you gonna write a function called python_function(numbers: List[float], threshold: float) >->- > bool that gon’ do this following functionality: Aight, Listen. Say you got a list of numbers yeah? Now, we trynna see if any two of ’em numbers is closer to each other than a number you give, feel me?So, this is what we ’bout to do: $>$much-greater-thanabsent>>>> > > python_function([1.0, 2.0, 3.0], 0.5) False That’s gon’ give you False cuz ain’t none of ’em numbers close enough.But, if you hit it like: $>$much-greater-thanabsent>>>> > > python_function([1.0, 2.8, 3.0, 4.0, 5.0, 2.0], 0.3) True Bet you gettin’ True, cuz this time some of ’em numbers real tight. You gotta whip up a Python function to handle this problem. You gon’ make sure the function name right, which gotta python_function.
Math
[…]
Math
[…]
Logic
[…]
Logic
[…]
Comprehensive
Standardized To try fishing for the first time, here are the steps and the times needed for each step Step 1. drive to the outdoor store (10 minutes) Step 2.compare fishing poles (30 minutes) Step 3. buy a fishing pole (5 minutes) Step 4. buy some bait (5 minutes) Step 5. drive to a lake (20 minutes) Step 6. rent a small boat (15 minutes)
These ordering constraints need to be obeyed when executing above steps: Step 1 must precede step 2. Step 2 must precede step 3. Step 2 must precede step 4. Step 3 must precede step 5. Step 4 must precede step 5 Step 5 must precede step 6.
[…]
Comprehensive
AAVE If you finna go fish for the first time, here’s what you got to know and the times you need for each step. Step 1. To kick things off, pull up to the outdoor store (10 minutes) Step 2. Check out which one of them fishing poles is good and which one is not (30 minutes) Step 3. Cop a fishing pole (5 minutes) Step 4.Get yourself some bait as well (5 minutes) Step 5. Head out to a lake (20 minutes) Step 6.rent yourself a small boat (15 minutes)
These ordering constraints gotta be followed when you doin’ ’em steps above: You gotta deal with 1 before hittin’ the 2. You gotta deal with 2 before hittin’ the 3. You gotta deal with 2 before hittin’ the 4. You gotta deal with 3 before hittin’ the 5. You gotta deal with 4 before hittin’ the 5. You gotta deal with 5 before hittin’ the 6.
[…]
A.3 Rubrics
A.3.1 Employment Information
We work with data vendors to employ 13 annotators in total for our task. For algorithm instance annotation, we specifically hire annotators with computer science backgrounds. Annotators are self-identified as proficient speakers of African American Vernacular English. We do not pose any hard constraints in verifying dialect identity as previous studies do (e.g., Ziems et al. (2023)). We note even within a dialect there can be significant variations on the individual level and that we want to avoid homogenization and over-simplification of the dialect (King, 2020). Instead, we ask self-identified annotators to cross-check each other’s annotations and modify if they sound unnatural.
Details of employment are shown below.
Information Collected We do not collect personally identifiable information from our annotators (e.g., name, age, etc). We only collect the annotators’ responses to our consent form and their annotations of our data.
Risk and Consent We note that our base datasets are from publicly available, widely used, peer-reviewed datasets that adhere to peer-review regulations. Moreover, our tasks are mainly centered around reasoning, which does not concern sensitive information per se. In addition, we make sure that annotators understand the risks of the annotation (i.e., although we have tried our best to ensure the safety of the data, it is still possible that they may feel uncomfortable in the annotation) and their right to exit the task during the process by signing a consent form prior to the start of the task.
Compensation We offer payment to annotators with hourly rates higher than the U.S. federal minimum wage.
No AI Assistant We explicitly inform our annotators that they should not reply on any AI assistant tools to help them complete the task. To further ensure this, we design our annotation platform to disallow copy and paste. The default annotation area for annotators is the original text, which means that it is easier for annotators to simply edit the text than querying AI assistants.
A.3.2 Annotation Guideline
You need to translate/rephrase/localize the task input in a way that is natural to the speakers of your dialect without changing the intention of the prompts. You should not change named entities, numbers, equations, variable names and other formal devices that are not natural language per se or those that would affect the intention of the prompts. The translation does not need to be grammatical or acceptable in standard English. Rather, it should accurately reflect the features of their dialects. You can add or delete some functional content to make the prompts sound more natural (e.g., adding fillers). However, you should keep the vital information complete and unchanged.
You should NOT change information that would invalidate the output given the question. If you are unsure about any specific parts, leave them unchanged. Especially, you should not change the following parts:
(i) numbers (e.g. 180 in 180 days)
(ii) units (e.g. days in 180 days)
[…]
(iv) proper nouns (e.g., Natalia in Natalia sold clips to 48 of her friends)
(v) function names, variables, data types, and input-output examples (e.g., $>$much-greater-thanabsent>>>> > > has_close_elements($[1.0,2.0,3.0]$, 0.5) False $>$much-greater-thanabsent>>>> > > has_close_elements($[1.0,2.8,3.0,4.0,5.0,2.0]$, 0.3) True in Check if in given list of numbers, are any two numbers closer to each other than given threshold. $>$much-greater-thanabsent>>>> > > has_close_elements($[1.0,2.0,3.0]$, 0.5) False $>$much-greater-thanabsent>>>> > > has_close_elements($[1.0,2.8,3.0,4.0,5.0,2.0]$, 0.3) True)
A.4 Data Quality Verification
After we conduct human validations for naturalness and correctness of prompts, we conduct the final round sanity check with GPT-4o. We prompt GPT-4o with temperature 0.7 and sample three instances for each query. We manually inspect instances again where all of the answers suggest that they are invalid paraphrases of the original prompts.
User prompt
You will be given two prompts, one in Standard English and one in African American English. Determine whether the African American English prompt is a valid paraphrase of the Standard English prompt. Ignore the semantic validaty of the Standard English prompt. Standard English: "[SAE_PROMPT]" African American English: "[AAVE_PROMPT]" Is the African American English prompt a valid paraphrase of the Standard English prompt?
A.5 Implementation Details
A.5.1 Dataset Implementation
For Algorithm, we unify the prompts by substituting all function names as python_function to avoid as much memorization as possible. We also manually corrected instances in HumanEval where the task descriptions were not precise enough (e.g., when the output data structure specified in the docstring is different from the one specified in the function heading). We also slightly modified some instructions in algorithm datasets without changing their intention to make sure our prompts are coherent (e.g., changing to solve the following problem to to realize the following functionality).
For other tasks, we unify the task output by asking LLMs to encode answers in $<answer></answer>$ to enable easy parsing. All details can be found in ReDial dataset files.
A.5.2 Inference Implementation
We set temperature=0 and max new token as 4096 for all models at inference time unless specified in the main paper. We run experiments on GPT-4o/4/3.5 via Azure OpenAI service. We evaluate all other models via Azure Machine Learning Studio API for main results. Experiments run in the analysis part are hosted on 4 A100 with 80GB memory each.
A.6 Results for Non-zero Temperature
We vary the temperature by 0, 0.5, 0.7, and 1 on GPT-4o/4/3.5-turbo and Phi-3-Mini/Medium-128K-Instruct. When the temperature is not 0, we sample 3 answers per query and take average pass rates as results for corresponding settings. Results are in Figure 3.
Figure 3: We vary the temperature by 0, 0.5, 0.7, 1 and report the performance gap between Standardized and AAVE ReDial.
We find that increasing temperature reduces the gap for GPT-4o in general, but does not affect other models’ performance as much. Even when the performance gap is reduced, increasing temperature cannot cancel the gap.
A.7 Full Results on ReDial
we present the complete results on Redial. Specifically, Table 5 provides the detailed results for Algorithm, Table 6 covers the results for Logic, Table 7 reports the results for Math, and Table 8 reports the results for Integrated Tasks.
Model Setting HumanEval MBPP
Original AAVE Original AAVE
[…]
CoT 0.841 0.805(-)0.037 0.693 0.713(+)0.02
[…]
CoT 0.750 0.707(-)0.043 0.693 0.500(-)0.193
[…]
CoT 0.616 0.591(-)0.024 0.680 0.507(-)0.173
[…]
CoT 0.793 0.726(-)0.067 0.753 0.747(-)0.007
LLaMA-3.1-70B-Instruct Vanilla 0.744 0.726(-)0.018 0.707 0.573(-)0.133
CoT 0.738 0.689(-)0.049 0.707 0.613(-)0.093
LLaMA-3-70B-Instruct Vanilla 0.689 0.671(-)0.018 0.673 0.613(-)0.06
CoT 0.720 0.665(-)0.055 0.673 0.627(-)0.047
LLaMA-3-8B-Instruct Vanilla 0.530 0.524(-)0.006 0.540 0.493(-)0.047
CoT 0.537 0.512(-)0.024 0.527 0.440(-)0.087
Mixtral-8x7B-Instruct-v0.1 Vanilla 0.402 0.390(-)0.012 0.507 0.413(-)0.093
CoT 0.396 0.396(-)-0.0 0.547 0.427(-)0.12
Mistral-7B-Instruct-v0.3 Vanilla 0.268 0.268(-)-0.0 0.400 0.240(-)0.16
CoT 0.262 0.274(+)0.012 0.367 0.213(-)0.153
Phi-3-Medium-128K-Instruct Vanilla 0.530 0.518(-)0.012 0.560 0.340(-)0.22
CoT 0.530 0.573(+)0.043 0.567 0.327(-)0.24
Phi-3-Small-128K-Instruct Vanilla 0.598 0.329(-)0.268 0.633 0.167(-)0.467
CoT 0.585 0.293(-)0.293 0.553 0.087(-)0.467
Phi-3-Mini-128K-Instruct Vanilla 0.549 0.482(-)0.067 0.567 0.367(-)0.2
CoT 0.567 0.530(-)0.037 0.587 0.347(-)0.24
Table 5: All results for Algorithm.
Model Setting Folio LogicBench
Original AAVE Original AAVE
[…]
CoT 0.938 0.926(-)0.012 0.715 0.645(-)0.070
[…]
CoT 0.864 0.759(-)0.105 0.735 0.730(-)0.005
[…]
CoT 0.519 0.506(-)0.012 0.490 0.360(-)0.130
[…]
CoT 0.907 0.877(-)0.031 0.820 0.730(-)0.090
LLaMA-3.1-70B-Instruct Vanilla 0.642 0.593(-)0.049 0.750 0.660(-)0.090
CoT 0.870 0.827(-)0.043 0.760 0.720(-)0.040
LLaMA-3-70B-Instruct Vanilla 0.673 0.623(-)0.049 0.655 0.495(-)0.160
CoT 0.883 0.809(-)0.074 0.400 0.360(-)0.040
LLaMA-3-8B-Instruct Vanilla 0.667 0.617(-)0.049 0.325 0.340(+)0.015
CoT 0.599 0.660(+)0.062 0.375 0.355(-)0.020
Mixtral-8x7B-Instruct-v0.1 Vanilla 0.327 0.401(+)0.074 0.485 0.110(-)0.375
CoT 0.370 0.284(-)0.086 0.395 0.285(-)0.110
Mistral-7B-Instruct-v0.3 Vanilla 0.481 0.537(+)0.056 0.180 0.055(-)0.125
CoT 0.475 0.506(+)0.031 0.200 0.120(-)0.080
Phi-3-Medium-128K-Instruct Vanilla 0.543 0.568(+)0.025 0.465 0.390(-)0.075
CoT 0.698 0.574(-)0.123 0.325 0.330(+)0.005
Phi-3-Small-128K-Instruct Vanilla 0.580 0.531(-)0.049 0.490 0.520(+)0.030
CoT 0.728 0.568(-)0.160 0.395 0.485(+)0.090
Phi-3-Mini-128K-Instruct Vanilla 0.420 0.352(-)0.068 0.755 0.665(-)0.090
CoT 0.481 0.370(-)0.111 0.735 0.655(-)0.080
Table 6: All results for Logic.
Model Setting GSM8K SVAMP
Original AAVE Original AAVE
[…]
CoT 0.967 0.933(-)0.033 0.933 0.907(-)0.027
[…]
CoT 0.947 0.867(-)0.080 0.893 0.760(-)0.133
[…]
CoT 0.780 0.480(-)0.300 0.727 0.607(-)0.120
[…]
CoT 0.973 0.960(-)0.013 0.933 0.920(-)0.013
LLaMA-3.1-70B-Instruct Vanilla 0.680 0.920(+)0.240 0.853 0.867(+)0.013
CoT 0.867 0.927(+)0.060 0.893 0.813(-)0.080
LLaMA-3-70B-Instruct Vanilla 0.933 0.920(-)0.013 0.880 0.853(-)0.027
CoT 0.947 0.907(-)0.040 0.900 0.867(-)0.033
LLaMA-3-8B-Instruct Vanilla 0.847 0.800(-)0.047 0.807 0.800(-)0.007
CoT 0.820 0.800(-)0.020 0.833 0.800(-)0.033
Mixtral-8x7B-Instruct-v0.1 Vanilla 0.427 0.193(-)0.233 0.613 0.487(-)0.127
CoT 0.673 0.573(-)0.100 0.700 0.560(-)0.140
Mistral-7B-Instruct-v0.3 Vanilla 0.367 0.147(-)0.220 0.433 0.280(-)0.153
CoT 0.420 0.320(-)0.100 0.487 0.373(-)0.113
Phi-3-Medium-128K-Instruct Vanilla 0.893 0.833(-)0.060 0.840 0.747(-)0.093
CoT 0.893 0.853(-)0.040 0.827 0.800(-)0.027
Phi-3-Small-128K-Instruct Vanilla 0.840 0.793(-)0.047 0.800 0.727(-)0.073
CoT 0.880 0.873(-)0.007 0.907 0.813(-)0.093
Phi-3-Mini-128K-Instruct Vanilla 0.520 0.573(+)0.053 0.520 0.527(+)0.007
CoT 0.800 0.807(+)0.007 0.747 0.693(-)0.053
Table 7: All results for Math.
Model Setting Original AAVE
[…]
CoT 0.762 0.662(-)0.1
[…]
CoT 0.283 0.058(-)0.225
[…]
CoT 0.075 0.067(-)0.008
[…]
CoT 0.900 0.771(-)0.129
LLaMA-3.1-70B-Instruct Vanilla 0.392 0.113(-)0.279
CoT 0.579 0.500(-)0.079
LLaMA-3-70B-Instruct Vanilla 0.158 0.067(-)0.092
CoT 0.517 0.350(-)0.167
LLaMA-3-8B-Instruct Vanilla 0.025 0.067(+)0.042
CoT 0.029 0.025(-)0.004
Mixtral-8x7B-Instruct-v0.1 Vanilla 0.100 0.075(-)0.025
CoT 0.133 0.071(-)0.062
Mistral-7B-Instruct-v0.3 Vanilla 0.096 0.075(-)0.021
CoT 0.083 0.083(-)-0.0
Phi-3-Medium-128K-Instruct Vanilla 0.050 0.037(-)0.013
CoT 0.067 0.029(-)0.037
Phi-3-Small-128K-Instruct Vanilla 0.058 0.062(+)0.004
CoT 0.096 0.079(-)0.017
Phi-3-Mini-128K-Instruct Vanilla 0.021 0.042(+)0.021
CoT 0.017 0.021(+)0.004
Table 8: All results for Integrated.
A.8 Perplexity of Typos vs. Natural AAVE Text
We gradually inject typos in SE ReDial by replacing/deleting/adding words/characters, such that we make the input texts more difficult for LLMs (i.e., the measured perplexity goes up). We find that while these perturbations degrade model performance, the drop did not reach the severity observed with natural AAVE data on large-scale models (see full results in Section A.8). This discrepancy suggests that AAVE brittleness is not solely due to the general difficulty in processing. Interestingly, larger models such as LLaMA-3.1-70B-Instruct and Phi-3-Medium-128K-Instruct degrade less on synthetic noise yet still struggle disproportionately with authentic AAVE data. This suggests their brittleness stems from reliance on SE priors and limited encoding of AAVE-specific discourse structures.
Figure 4: Model performance on misspelled SE compared to human-written AAVE data. We gradually add noise to SE ReDial to increase its perplexities until they surpass the perplexity of AAVE ReDial and report the models’ performance on every perturbation level. Horizontal and vertical lines refer to model pass rates/perplexities on AAVE ReDial respectively. Larger LLMs (i.e., LLaMA-3.1-70B-Instruct and Phi-3-Medium-128K-Instruct) perform worse on AAVE than on perturbed text with a similar perplexity level.
A.9 Multivalue Perturbation
Since the unfamiliarity of data cannot explain the whole picture, how much can we attribute the failure to AAVE-specific features? We use the rule-based transformation method in Ziems et al. (2023) to inject AAVE features into our dataset for synthetic probing. We compare GPT-4o/4/3.5 and Phi-3-Medium/Mini-128k-Instruct performance in feature densities of $\{0,0.25,0.5,0.75,1\}$ and run the same setting as the main experiment.
Figure 5: Perturbation with AAVE features. We control perturbation feature densities at $\{0,0.25,0.5,0.75,1\}$ to gradually inject AAVE features using rule-based transformations.
Results are shown in Figure 5. On the one hand, we find that models generally show increasing performance drops with increasing feature density, which means that AAVE-specific features do contribute to model performance drops. On the other hand, even drops caused by the strongest perturbation are generally far from the drops caused by human-rewritten prompts. This shows the limitation of previous methods in revealing LLM robustness based on synthetic data as there can be more influential factors than what lexico-syntactic rules can capture. Phi-3-Mini-128K-Instruct is again an outlier here, being that it is the only model that has a stronger performance drop in feature injections compared to human-written dialect data.
A.10 Qualitative Analysis of Standardized Prompting
We observe a sensible gap between model performance on Standardized ReDial with zero-shot prompting and AAVE ReDial with standardization prompting. In this section, we qualitatively compare GPT-4o’s outputs in these two settings, the model with the best absolute overall performance, and examine its errors. We focus on the math subset of ReDial and identify three key error patterns: wrong question rephrasing, distraction by irrelevant information, and failure to execute all steps.
Wrong question rephrasing. GPT-4o wrongly phrases question ‘Jame … How many years have they got between them now if in 8 years his cousin will be 5 years younger than twice his age?’ to ‘James … How old is his cousin now?’, which changes the question of age gap to absolute age.
Distraction by irrelevant information. GPT-4o gets distracted by task-irrelevant information after AAVE standardization while the distraction is not observed in Standardized ReDial. For instance, in ‘Say we got 8 different books and 10 different movies in the crazy silly school series. How many more movies than books is there gon be in the crazy silly school series if you read 19 books and watched 61 movies?’, books that have been read and movies that have been watched are not associated with the answer. Although GPT-4o can ignore irrelevant information in Standardized ReDial, it gets distracted after AAVE standardization, which shows the brittleness of its reasoning ability.
Failure to execute all the steps. GPT-4o sometimes simulates an algorithm to solve math problems after standardization (e.g., ‘Let $(x)$ be the number of apple pie boxes…’). However, it does not fully solve the problem in the end and only returns a formula (e.g., ‘30x + 255’), which indicates that the model’s reasoning ability is limited when it comes to program simulation for queries expressed in dialects.
A.11 Synthetic Lexical Feature list
Following Ziems et al. (2022), we implement a feature list with distinct AAVE lexicons: [’got’, ’ain’t’, ’no’, "’", ’gonna’, ’wanna’, ’gotta’, ’done’, ’been’, ’finna’, ’gunna’, ’gon’,], and compute the Spearman’s correlation between their frequencies in AAVE ReDial and the performance drop.
A.12 Statement of Contribution
All co-authors contributed to discussions, provided input on various aspects of the project, and assisted with writing, editing, and advising. In addition to these contributions, FL developed the initial idea, designed and conducted the experiments, contributed significantly to data collection, drafted the paper, and performed all analyses unless otherwise specified. As FL’s mentors during her internship at Microsoft, SM and AW contributed significantly by coordinating resources, guiding the overall direction of the project including data collection, managing the ethical review process, and serving as the primary corresponding authors. XW conducted experiments on Claude and GPT-o1. On top of advising and paper writing, ELM developed some of the initial experiments for the ablation study and VH contributed to ideation and experiment design.